# Optimization of modulation transfer protocol for Rydberg RF receivers

### v1

@@ Abstract
We explore theoretically and experimentally the recently demonstrated modulation transfer protocol [D.-A. Trinh, K. V. Adwaith, M. Branco, A. Rouxel, S. Welinski, P. Berger, F. Goldfarb, and F. Bretenaker, Applied Physics Letters 125, 154001 (2024)] aiming at extending the bandwidth of quantum RF receivers based on hot Rydberg atoms. This protocol is based on a phase modulation of the coupling beam, which is transformed by the nonlinear response of the atoms into an amplitude modulation of the probe beam. We develop a theoretical model to optimize both the modulation frequency and the modulation amplitude of the coupling beam, thereby maximizing the atomic response. Once optimized, the sensitivity to detuned RF fields of this modulation transfer protocol is compared with that of the conventional protocol. This comparison shows that the new protocol outperforms the usual one as soon as the RF signal to be measured is detuned by more than a few MHz and offers a complementary approach to increase the detection bandwidth. In all cases, the experimental results are in good agreement with the simulations.
@@ I Introduction
A large fraction of quantum sensors are based on atomic resonances probed by electromagnetic radiation. In such cases, the sensor sensitivity typically depends on i) the strength of the probed resonance and ii) the linewidth of the resonance [budkerOpticalMagnetometry2007]. According to these two criteria, Rydberg atoms, i.e. atomic levels corresponding to large values of the principal quantum number $n$, are particularly appealing because i) the electric dipole moment between two neighboring Rydberg levels scales like $n^{2}$ and ii) the lifetime of Rydberg levels can be as long as milliseconds, making resonances between them extremely narrow.
Since the transitions between Rydberg levels exhibit frequencies ranging from the MHz to the THz domain, these considerations have led to the development of very promising RF receivers based on such Rydberg levels. These RF receivers rely on a three-level Electromagnetically Induced Transparency (EIT) [finkelsteinPracticalGuideElectromagnetically2023] in a thermal vapor of alkali atoms, where the RF field induces an Autler-Townes (AT) splitting. This technique was initially introduced using Rubidium-87 ($\mathrm{{}^{87}Rb}$) [sedlacekMicrowaveElectrometryRydberg2012] and has since been expanded to include Rubidium-85 ($\mathrm{{}^{85}Rb}$) [hollowayBroadbandRydbergAtomBased2014] and Cesium [fanDispersiveRadioFrequency2016].
While the primary focus of such sensors is electrometry through the measurement of RF field amplitude, some studies have developed alternative methods using the same initial apparatus. These methods aim to measure the phase [simonsRydbergAtombasedMixer2019, andersonOpticalRadioFrequencyPhase2022, jingAtomicSuperheterodyneReceiver2020], the polarization [sedlacekAtomBasedVectorMicrowave2013, wangPreciseMeasurementMicrowave2023], or even the angle of arrival of the RF field [robinsonDeterminingAngleofarrivalRadiofrequency2021, yanThreedimensionalLocationSystem2023].
This all-optical detection method allows for detecting signals ranging from the sub-Megahertz [jauVaporCellBasedAtomicElectrometry2020, liuHighlySensitiveMeasurement2022, limKilohertzrangeElectricField2023, brownVeryhighUltrahighfrequencyElectricfield2023] range to Terahertz frequencies [chenTerahertzElectrometryInfrared2022], exploiting the large number of allowed atomic transitions between Rydberg states in alkali atoms. These transitions exhibit large electric dipole moments, which make Rydberg RF receivers highly sensitive to resonant microwave fields.
@@
To address this limitation, some research groups have proposed to improve the RF bandwidth by inducing an AC Stark shift using a high-intensity [andersonContinuousfrequencyMeasurementsHighintensity2017] detuned RF field. This approach can create a large bandwidth of several Gigahertz (ranging from 2 to 5 GHz), but results in a reduced sensitivity, which is not ideal for detecting weak RF fields. Another approach consists in using an RF local oscillator [huContinuouslyTunableRadio2022, yaoSensitivityEnhancementFardetuned2022, liRydbergAtombasedAM2022] to enhance the sensitivity and enable detection of weak RF fields. It has also been demonstrated that exciting a second Rydberg transition permits to tune the sensor resonance frequency [simonsContinuousRadiofrequencyElectricfield2021]. However, both of these techniques require a second RF field, which undermines the goal of developing an all-optical detection system. Alternatively, DC Stark shifts [ouyangContinuousBroadbandMicrowave2023] can be used, but they necessitate electrodes inside the vapor cell, which also contradicts the aim of creating an all-dielectric detection system.
In order to overcome these limitations, an all-optical method has been demonstrated to increase the RF bandwidth using a four-wave mixing modulation transfer [heNoiseSpectroscopyRydberg2021, trinhModulationTransferProtocol2024]. It is based on the use of a phase modulation of the coupling laser, which is transferred to the probe laser through the nonlinearity of the atomic medium. A new signal is then obtained by detecting the intensity modulation of the probe field, which has been shown to allow a significant improvement of the RF detection bandwidth [trinhModulationTransferProtocol2024].
The aim of this article is thus to optimize this modulation transfer protocol (MTP) in order to expand the RF bandwidth as much as possible compared to the conventional protocol (CP). This is performed by optimizing the modulation parameters in a three-level ladder EIT scheme in $\mathrm{{}^{85}Rb}$. We begin by developing a theoretical model to simulate the response of the system along the CP and the MTP. This model is subsequently used to optimize the modulation parameters. The results from these simulations are then compared to experimental measurements. Next, we analyze the response of the MTP to weak RF fields and compare it to the response of the CP. Finally, we investigate the improvement of the RF bandwidth.
@@ II Theory and simulations
As stated in the introduction, this section is dedicated to the theory and simulation of the Rydberg sensor along the two protocols. In the conventional protocol (CP, Section II.A), variations in DC transmission of the probe field are used to measure the RF field. In the modulation transfer protocol [trinhModulationTransferProtocol2024], the coupling beam is phase modulated leading to a modulation of the probe beam intensity (Section II.B).)
@@ II.1 Conventional protocol
$\mathcal{E}_{\mathrm{p}}(x+dx)=\mathcal{E}_{\mathrm{p}}(x)-\frac{\mathrm{i}\omega_{\mathrm{p}}}{2\varepsilon_{0}c}\mathcal{N}_{0}\wp_{12}dx\int_{-\infty}^{+\infty}\rho_{21}(x,v)P(v)dv\ ,$ (4)
@@
In the following, we focus on the following levels of 85Rb: level $\left|1\right$> is $5^{2}S_{1/2}(F=3)$, level $\left|2\right$> is $5^{2}P_{3/2}(F=4)$, level $\left|3\right$> is $50^{2}D_{5/2}(F=3,4,5)$, and level $\left|4\right$> corresponds to $51^{2}P_{3/2}(F=2,3,4)$. The values of parameters $\omega_{ij}$, $\wp_{ij}$, and $\Gamma_{i}$ are obtained from the Python package called Alkali Rydberg Calculator (ARC) [sibalicARCOpensourceLibrary2017]. The values of the elements of the electric dipole matrix are obtained by averaging over the transitions between the different sublevels that constitute each transition. The values of these parameters used in the simulations are compiled in Table 1. Moreover, the small signal transmission of the cell at resonance is equal to 0.34, and the atom transit rate in the beam is $\gamma_{\mathrm{t}}/2\pi=650\,\mathrm{kHz}$. All these values are deduced from the parameters of the experiment described below (see Section III.A).
@@
Population Rabi
Dipole ($ea_{0}$) Lifetime (MHz) Frequency (MHz)
@@ II.2 Modulation transfer protocol
The solution to this equation can be approximated by performing a Floquet expansion of the density matrix, namely $\rho=\sum_{n}\rho^{(n)}\mathrm{e}^{-\mathrm{i}\omega_{\mathrm{mod}}t}$. The amplitudes $\rho^{(n)}$ of the different frequency components can then be obtained using the continuous fraction method [wongInfluenceCoherentRaman2004]. By truncating these continuous fractions at orders $\pm 1$, one obtains the following system of equations:
@@
Equation (8) shows that the atoms radiate light at frequencies $\omega_{\mathrm{p}}$ and $\omega_{\mathrm{p}}\pm\omega_{\mathrm{mod}}$. This transfer of modulation from the coupling to the probe beam, which can be interpreted as a four-wave mixing phenomenon mediated by the nonlinearity of the atoms [ducloyTheoryDegenerateFourwave1982], adds sidebands to the probe beam. This induces a modulation of the intensity of this beam, as depicted in Fig.1(d). We can thus extract from the probe beam detection a novel signal known as the Relative Modulation Amplitude (R.M.A), defined as:
@@ III.1 Experimental apparatus
The probe and coupling lasers at 780 and 480 nm, respectively, propagate in opposite directions through a quartz cell, which contains a natural mixture of Rubidium isotopes at room temperature. Within the cell, the probe (coupling) laser has a waist diameter of 300 (400) $\mathrm{\mu m}$, leading to Rayleigh ranges longer than the 7.5 cm length of the vapor cell. The input laser powers are $0.4\,\mu\mathrm{W}$ for the probe laser and $46\,\mathrm{mW}$ for the coupling laser. A horn antenna (MVG QR18000) located at 57.5 cm from the vapor cell generates the RF field, which is sent to the cell orthogonally to the laser beams. To shield the antenna and vapor cell from external fields, we use a custom-built small anechoic chamber made of aluminum sheets and lined with RF-absorbing panels. The probe, coupling, and RF fields are all linearly polarized along the Y-axis.
@@
The frequency of the 780 nm probe laser (Toptica DL-Pro) is locked to the transition between levels $5^{2}S_{1/2}(F=3)$ and $5^{2}P_{3/2}(F=4)$ of 85Rb using modulation transfer spectroscopy [mccarronModulationTransferSpectroscopy2008, sunModulationTransferSpectroscopy2016, preuschoffOptimizationStrategiesModulation2018, limModulationTransferSpectroscopy2022] in an auxiliary cell. The 480 nm coupling laser (Toptica TA-SHG) is locked to the transition between the $5^{2}P_{3/2}(F=4)$ and $50^{2}D_{5/2}(F=3,4,5)$ levels using an EIT resonance in the presence of the probe laser [abelLaserFrequencyStabilization2009]. The amplitude of the RF electric field $E_{\mathrm{RF}}$ at the position of the vapor cell is estimated after accounting for the losses, the gain of the antenna, and its distance from the vapor cell. The cell perturbation factor [robinsonAtomicSpectraSixlevel2021, fanEffectVaporCellGeometry2015, rotunnoInvestigatingElectromagneticallyInduced2023], which reduces the average RF field within the cell, is estimated to be approximately 0.54 according to experimental characterization using AT splitting. This RF signal couples level $50^{2}D_{5/2}(F=3,4,5)$ with level $51^{2}P_{3/2}(F=2,3,4)$.
@@
To tune its frequency, the coupling beam passes through two single-pass AOMs inducing frequency shifts of opposite signs. The second AOM is also used to phase modulate [liOpticalPhaseShifting2005] the coupling beam in the modulation transfer protocol. The phase modulation frequency $\omega_{\mathrm{mod}}$ and amplitude $\beta$ are programmed using an arbitrary waveform generator. The value of $\beta$ is experimentally calibrated by measuring the laser spectrum with a Fabry-Perot interferometer. In the modulation transfer protocol, the probe beam intensity is demodulated at $\omega_{\mathrm{mod}}$, and the two quadratures provided by this demodulation are combined to measure the R.M.A defined in Eq. (9). In the conventional protocol, the coupling beam intensity is modulated by a square signal at 100 kHz with a 100 % modulation depth, and the probe beam intensity is demodulated with a lock-in detection. This directly provides the so-called probe beam transparency, which is the difference in probe beam transmission when the coupling beam is present and when it is absent. Similarly to the coupling beam, the probe beam frequency can also be scanned thanks to two acousto-optic modulators. In all cases, the probe beam is detected with a fast avalanche photodiode (APD) (Thorlabs APD410A/M) with a 10 MHz bandwidth and $1.2\times 10^{7}\ \mathrm{V/W}$ sensitivity at 780 nm. Another photodetector detects a fraction of the incident probe power on the cell to calibrate transmission measurements.
@@ III.2 Optimization of the Phase Modulation of the Coupling Beam
Now that we have checked the validity of our simulations, we can use them to optimize the response of the modulation transfer protocol. In particular, we have shown in Ref. [trinhModulationTransferProtocol2024] that the modulation transfer protocol is particularly interesting to enhance the bandwidth of RF detection. This improvement comes from the sharp dip that can be observed around $\Delta_{\mathrm{p}}/2\pi=0$ in Figs. 2 and 4. This dip, which originates from destructive interference between the two sidebands of the coupling beam, creates a steep slope near the probe resonance $\Delta_{\mathrm{p}}/2\pi=0\,\mathrm{MHz}$. When the RF field is detuned, one can use this steep slope at $\Delta_{\mathrm{p}}/2\pi$ slightly different from 0 (typically $100\,\mathrm{kHz}$), to enhance the response of the R.M.A signal to the variations of $E_{\mathrm{RF}}$.
@@
The above discussion, based on the results of Fig. 6, confirms that the optimized MTP permits to extend the sensitivity of the sensor to detuned RF fields. In the following section, we provide a quantitative comparison of the sensitivity of the two protocols and of the bandwidth improvement provided by the optimized MTP.

### v2

# Bridging gaps in Rydberg RF receivers using modulation transfer bandwidth enhancement

@@ Abstract
We optimize theoretically and experimentally the performances of the recently demonstrated modulation transfer protocol [D.-A. Trinh, K. V. Adwaith, M. Branco, A. Rouxel, S. Welinski, P. Berger, F. Goldfarb, and F. Bretenaker, Applied Physics Letters 125, 154001 (2024)] aiming at extending the bandwidth of quantum RF receivers based on hot Rydberg atoms. This optimization relies on tuning the parameters of the phase modulation of the coupling beam, which is converted by the nonlinear response of the atoms into an amplitude modulation of the probe beam. We develop a theoretical model to optimize both the modulation frequency and the modulation amplitude of the coupling beam, thereby maximizing the atomic response. Once optimized, the sensitivity to detuned RF fields of this modulation transfer protocol is measured and compared with that of the conventional protocol. This comparison shows that the new protocol permits a strong increase in the detection bandwidth. Indeed, it outperforms the usual one as soon as the RF signal to be measured is detuned by more than a few MHz. We illustrate the capability of this modulation transfer protocol to enhance the detection bandwidth by showing experimentally how it permits to bridge the gap between two Rydberg transitions separated by 166 MHz. In all cases, the experimental results are in good agreement with the simulations.
@@ I Introduction
A large fraction of quantum sensors are based on atomic resonances probed by electromagnetic radiation. In such cases, the sensor sensitivity typically depends on i) the strength of the probed resonance and ii) the linewidth of the resonance [5]. According to these two criteria, Rydberg atoms, i.e. atomic levels corresponding to large values of the principal quantum number $n$, are particularly appealing because i) the electric dipole moment between two neighboring Rydberg levels scales like $n^{2}$ and ii) the lifetime of Rydberg levels can be as long as milliseconds, making resonances between them extremely narrow.
Since the transitions between Rydberg levels exhibit frequencies ranging from the MHz to the THz domain, these considerations have led to the development of very promising RF receivers based on such Rydberg levels. These RF receivers rely on a three-level Electromagnetically Induced Transparency (EIT) [11] in a thermal vapor of alkali atoms, where the RF field induces an Autler-Townes (AT) splitting. This technique was initially introduced using Rubidium-87 ($\mathrm{{}^{87}Rb}$) [31] and has since been expanded to include Rubidium-85 ($\mathrm{{}^{85}Rb}$) [13] and Cesium [8].
While the primary focus of such sensors is electrometry through the measurement of RF field amplitude, some studies have developed alternative methods using the same initial apparatus. These methods aim to measure the phase [34, 3, 16], the polarization [30, 37], or even the angle of arrival of the RF field [27, 39].
This all-optical detection method allows for detecting signals ranging from the sub-Megahertz [15, 22, 20, 4] range to Terahertz frequencies [6], exploiting the large number of allowed atomic transitions between Rydberg states in alkali atoms. These transitions exhibit large electric dipole moments, which make Rydberg RF receivers highly sensitive to resonant microwave fields.
@@
To address this limitation, some research groups have proposed to improve the RF bandwidth by inducing an AC Stark shift using a high-intensity [2] detuned RF field. This approach can create a large bandwidth of several Gigahertz (ranging from 2 to 5 GHz), but results in a reduced sensitivity, which is not ideal for detecting weak RF fields. Another approach consists in using an RF local oscillator [14, 40, 19] to enhance the sensitivity and enable detection of weak RF fields. It has also been demonstrated that exciting a second Rydberg transition permits to tune the sensor resonance frequency [33]. However, both of these techniques require a second RF field, which undermines the goal of developing an all-optical detection system. Alternatively, DC Stark shifts [24] can be used, but they necessitate electrodes inside the vapor cell, which also contradicts the aim of creating an all-dielectric detection system.
In order to overcome these limitations, an all-optical method has been demonstrated to increase the RF bandwidth by using a modulation transfer similar to four-wave mixing [12, 36]. It is based on the use of a phase modulation of the coupling laser, which is transferred to the probe laser through the nonlinearity of the atomic medium. A new signal is then obtained by detecting the intensity modulation of the probe field, which has been shown to allow a significant improvement of the RF detection bandwidth [36].
The aim of this article is thus to improve this modulation transfer protocol (MTP) to expand the RF bandwidth as much as possible compared to the conventional protocol (CP), thereby closing the gap between two closely spaced RF transitions. This is performed by optimizing the modulation parameters in a three-level ladder EIT scheme in $\mathrm{{}^{85}Rb}$. The first step consists in developing and using a reliable theoretical model. This is performed in Section II, where we simulate the response of the system along the CP and the MTP. This model is subsequently used to optimize the modulation parameters. The results from these simulations are then compared with the experimental measurements in Section III. Then, Section IV provides a detailed comparison between the performances of the two protocols. The signal and the noise are carefully measured and analyzed to determine the conditions in which the MTP outperforms the usual approach. A complete comparison of the bandwidths and sensitivities is provided, and substantiated with simulations. Finally, the interest of the new approach is illustrated by showing how it permits maintaining a good sensitivity while the RF frequency is tuned between two adjacent Rydberg transitions, whereas the conventional approach leads to a complete loss of sensitivity.
@@ II Theory and simulations
As stated in the introduction, this section is dedicated to the theory and simulation of the Rydberg sensor along the two protocols. In the conventional protocol (CP, Section II.A), variations in DC transmission of the probe field are used to measure the RF field. In the modulation transfer protocol [36], the coupling beam is phase modulated leading to a modulation of the probe beam intensity (Section II.B). The simulations detailed in this section will be used in Section III.B to optimize the modulation parameters.
@@ II.1 Conventional Protocol
$\mathcal{E}_{\mathrm{p}}(x+dx)=\mathcal{E}_{\mathrm{p}}(x)+\frac{\mathrm{i}\omega_{\mathrm{p}}}{2\varepsilon_{0}c}\mathcal{N}_{0}\wp_{12}dx\int_{-\infty}^{+\infty}\rho_{21}(x,v)P(v)dv\ ,$ (4)
@@
In the following, we focus on the following levels of 85Rb: level $\left|1\right$> is $5^{2}S_{1/2}(F=3)$, level $\left|2\right$> is $5^{2}P_{3/2}(F=4)$, level $\left|3\right$> is $50^{2}D_{5/2}(F=3,4,5)$, and level $\left|4\right$> corresponds to $51^{2}P_{3/2}(F=2,3,4)$. The values of parameters $\omega_{ij}$, $\wp_{ij}$, and $\Gamma_{i}$ are obtained from the Python package called Alkali Rydberg Calculator (ARC) [32]. The values of the elements of the electric dipole matrix are obtained by averaging over the transitions between the different sublevels that constitute each transition. The values of these parameters used in the simulations are compiled in Table 1. Moreover, the small signal transmission of the cell at resonance is equal to 0.34, and the atom transit rate in the beam is $\gamma_{\mathrm{t}}/2\pi=650\,\mathrm{kHz}$. All these values are deduced from the parameters of the experiment described below (see Section III.A).
@@
Dipole Population Rabi
Moment ($ea_{0}$) Decay (MHz) Frequency (MHz)
@@ II.2 Modulation Transfer Protocol
The solution to this equation can be approximated by performing a Floquet expansion of the density matrix, namely $\rho=\sum_{n}\rho^{(n)}\mathrm{e}^{-\mathrm{i}\omega_{\mathrm{mod}}t}$. The amplitudes $\rho^{(n)}$ of the different frequency components can then be obtained using the continuous fraction method [38]. By truncating these continuous fractions at orders $\pm 1$, one obtains the following system of equations:
@@
Equation (8) shows that the atoms radiate light at frequencies $\omega_{\mathrm{p}}$ and $\omega_{\mathrm{p}}\pm\omega_{\mathrm{mod}}$. This transfer of modulation from the coupling to the probe beam, which can be interpreted as a four-wave mixing phenomenon mediated by the nonlinearity of the atoms [7], adds sidebands to the probe beam. For example, we see that the combination of the three incident frequencies $\omega_{\mathrm{c}}$, $\omega_{\mathrm{c}}\pm\omega_{\mathrm{mod}}$, and $\omega_{\mathrm{p}}$ creates the new side band frequency $\omega_{\mathrm{p}}\pm\omega_{\mathrm{mod}}$.
The appearance of these new sidebands induces a modulation of the intensity of the probe beam, as depicted in Fig. 1(d). We can thus extract from the probe beam detection a novel signal known as the Relative Modulation Amplitude (R.M.A), defined as:
@@ III.1 Experimental Apparatus
[…]
@@
The frequency of the 780 nm probe laser (Toptica DL-Pro) is locked to the transition between levels $5^{2}S_{1/2}(F=3)$ and $5^{2}P_{3/2}(F=4)$ of 85Rb using modulation transfer spectroscopy [23, 35, 25, 21] in an auxiliary cell. The 480 nm coupling laser (Toptica TA-SHG) is locked to the transition between the $5^{2}P_{3/2}(F=4)$ and $50^{2}D_{5/2}(F=3,4,5)$ levels using an EIT resonance in the presence of the probe laser [1]. The amplitude of the RF electric field $E_{\mathrm{RF}}$ at the position of the vapor cell is estimated after accounting for the losses, the gain of the antenna, and its distance from the vapor cell. The cell perturbation factor [26, 10, 28], which reduces the average RF field within the cell, is estimated to be approximately 0.54 according to experimental characterization using AT splitting. This RF signal couples level $50^{2}D_{5/2}(F=3,4,5)$ with level $51^{2}P_{3/2}(F=2,3,4)$.
@@
To tune its frequency, the coupling beam passes through two single-pass AOMs inducing frequency shifts of opposite signs. The second AOM is also used to phase modulate [18] the coupling beam in the modulation transfer protocol. The phase modulation frequency $\omega_{\mathrm{mod}}$ and amplitude $\beta$ are programmed using an arbitrary waveform generator. This allows us to fine-tune $\beta$ and to only provide first-order sidebands to the modulation. The value of $\beta$ is experimentally calibrated by measuring the laser spectrum with a Fabry-Perot interferometer. In the modulation transfer protocol, the probe beam intensity is demodulated at $\omega_{\mathrm{mod}}$, and the two quadratures provided by this demodulation are combined to measure the R.M.A defined in Eq. (9). In the conventional protocol, the coupling beam intensity is modulated by a square signal at 100 kHz with a 100 % modulation depth, and the probe beam intensity is demodulated with a lock-in detection. This directly provides the so-called probe beam transparency, which is the difference in probe beam transmission when the coupling beam is present and when it is absent. Similarly to the coupling beam, the probe beam frequency can also be scanned thanks to two acousto-optic modulators. In all cases, the probe beam is detected with a fast avalanche photodiode (APD) (Thorlabs APD410A/M) with a 10 MHz bandwidth and $1.2\times 10^{7}\ \mathrm{V/W}$ sensitivity at 780 nm. Another photodetector detects a fraction of the incident probe power on the cell to calibrate transmission measurements.
@@ III.2 Optimization of the Phase Modulation of the Coupling Beam
Now that we have checked the validity of our simulations, we can use them to optimize the response of the modulation transfer protocol. In particular, we have shown in Ref. [36] that the modulation transfer protocol is particularly interesting to enhance the bandwidth of RF detection. This improvement comes from the sharp dip that can be observed around $\Delta_{\mathrm{p}}/2\pi=0$ in Figs. 2 and 4. This dip, which originates from destructive interference between the two sidebands of the coupling beam, creates a steep slope near the probe resonance $\Delta_{\mathrm{p}}/2\pi=0\,\mathrm{MHz}$. When the RF field is detuned, one can use this steep slope at $\Delta_{\mathrm{p}}/2\pi$ slightly different from 0 (typically $100\,\mathrm{kHz}$), to enhance the response of the R.M.A signal to the variations of $E_{\mathrm{RF}}$.
@@
The above discussion, based on the results of Fig. 6, confirms that the optimized MTP permits to extend the sensitivity of the sensor to detuned RF fields. In the following section, we provide a quantitative comparison of the sensitivity of the two protocols and of the bandwidth improvement provided by the optimized MTP. We also demonstrate the MTP’s ability to close the gap between two nearby RF transitions.